# Generating and Understanding Tests via Path-Aware Symbolic Execution with LLMs

### v1

@@ V-B LLM Test Input Generation
Large Language Models (LLMs) have recently gained popularity in unit test generation. Many approaches focus on enhancing prompts with additional context or documentation. ChatUnitTest [19] parses the project to extract focal method information and dependencies to enrich the test prompt. Vikram et al. [20] leverage API documentation to generate property-based tests. Other works [21, 22] evaluate the impact of prompt design on code coverage and bug detection. HITS [23] employs program slicing to guide LLMs in generating tests incrementally. TELPA [24] enhances test generation using bidirectional method analysis and counterexample-guided refinement. IntUT [25] introduces explicit test intentions (e.g., inputs, mocks, and expected outputs) to guide LLMs, improving test quality but without exploring specific program paths.
@@
A growing body of work integrates LLMs with symbolic execution, although they lack fine-grained path visibility. LLMSym [26] uses LLMs to translate Python code into SMT formulas for test generation, but is limited to basic types and constrained by solver expressiveness. Cottontail [27] uses LLMs to solve SMT-formatted constraints for path satisfiability and also structured input validity. DeFiAligner [28] performs symbolic execution on smart contract bytecode and employs LLMs to check alignment between SMT-formatted symbolic summary and documentation. SymPrompt [29] performs LLM-based symbolic execution by embedding if-else branches in prompts, but cannot model loops or recursion. HyLLFuzz [30] calls LLMs when fuzzing stalls to generate path-targeted inputs. AutoEXE [31] replaces SMT solvers with LLM-based reasoning over code-based path constraints, but it does not decompose paths within loops or recursion, nor provide visibility into which paths are actually exercised.
In contrast, Palm explicitly enumerates program paths and constructs executable path-specific variants with interleaved assertions, enabling test validation and iterative refinement. More importantly, prior work lacks visual interfaces for comprehending LLM generated test behavior at the path level. None has been been evaluated through systematic user studies on path-aware reasoning of generated tests and disambiguation of redundant or missing tests in terms of path behavior.
VI Conclusion

### v2

@@ 6. Related Work
Large Language Models (LLMs) have recently gained popularity for unit test generation. Prior work mainly improves generation quality by enriching prompts or adding analysis-driven feedback. For example, ChatUniTest (Xie et al., 2023) augments prompts with focal-method context and dependencies, while Vikram et al. (Vikram et al., 2023) leverage API documentation to synthesize property-based tests. Other studies examine how prompt design affects coverage and bug finding (Guilherme and Vincenzi, 2023; Li and Doiron, 2023). More analysis-guided approaches include HITS (Wang et al., 2024b), which uses slicing to incrementally generate tests, and TELPA (Yang et al., 2024), which refines tests using program analysis and feedback from ineffective attempts. IntUT (Nan et al., 2025) introduces explicit test intentions to guide generation, but it does not aim to exercise specific control-flow paths.
@@
A growing line of work combines LLMs with symbolic execution or constraint reasoning. LLMSym (Wang et al., 2024a) uses LLMs to assist in translating code or constraints into SMT for test generation, while Cottontail (Tu et al., 2025) uses LLMs to help satisfy SMT-formatted constraints and structured-input validity. DeFiAligner (Deng et al., 2025) applies symbolic execution on smart-contract bytecode and uses LLMs to check consistency between symbolic summaries and documentation. SymPrompt (Ryan et al., 2024) embeds branch conditions into prompts to steer LLM reasoning, and HyLLFuzz (Meng et al., 2024) invokes LLMs to propose new inputs when fuzzing stalls. AutoEXE (Li et al., 2025) replaces SMT solving with LLM-based reasoning over code-level path constraints, but it provides limited transparency into which paths are actually exercised, especially in the presence of loops or recursion.
In comparison, PALM enumerates paths and constructs executable assertion-instrumented variants, enabling validated, iterative test generation with a visual interface to inspect path-level behavior and identify redundant or missing tests.
7. Conclusion
@@
Data-availability Statement
Our implementation and evaluation artifacts are available at https://github.com/UCLA-SEAL/PALM.
Acknowledgements.
This work is supported by the National Science Foundation under grant numbers 2426162, 2106838, 2106404 and 2106420. It is also supported in part by funding from Amazon and Samsung. We want to thank the anonymous reviewers for their constructive feedback that helped improve the work.